# A Concept-Centric Approach to Multi-Modality Learning

### v1

@@ 4.1 Pretraining
4.2 Image-Text Matching
Image-text matching is a binary classification task on whether a natural language sentence describes an image. Our framework can naturally adopt a common approach involving creating representations for sentences and images in a shared latent space. In contrast to those works, however, our latent space is a knowledge-embedded concept space that supports efficient probing. Specifically, given an image-text pair $(x_{m}^{\text{vision}},x_{n}^{\text{NL}})$, their representations in the learned concept space $\mathcal{C}$ are generated by $f_{\text{vision}}(x_{m}^{\text{vision}})=\Omega_{m}^{\text{vision}}$ and $f_{\text{NL}}(x_{n}^{\text{NL}})=\Omega_{n}^{\text{NL}}$. The probability that $(x_{m}^{\text{vision}},x_{n}^{\text{NL}})$ is a positive pair can be determined by the cross entailment probability of $P(\text{matched}|(x_{m}^{\text{vision}},x_{n}^{\text{NL}}))=\frac{1}{2}\left[P(\Omega_{m}^{\text{vision}}|\Omega_{n}^{\text{NL}})+P(\Omega_{n}^{\text{NL}}|\Omega_{m}^{\text{vision}})\right]$. This inference process is demonstrated in Fig. 7 in Appendix.
@@ 4.2 Image-Text Matching
To compare our framework’s performance, we implement other benchmark multi-modality models with applications in the Image-Text Matching task. The results are summarized in Table 2. In contrast to those models with traditional black-box architectures, our framework displays a more efficient learning process and adopts a more transparent inference process without sacrificing its performance. Details of this experiment can be found in Appendix D.
@@
4.3 Visual Question Answering
@@
Table 3: A comparison between our framework’s performance and state-of-the-art models. *indicates the method does not use program annotations.
Inference Process. A correctly predicted program sequence $\hat{z_{i}}$ starts with a Scene function that returns all objects in an image and ends with a program that outputs the answer $o_{i}$. Intermediate programs takes output from previous programs as inputs, which is a reoccurring process until the last function. Our concept space $\mathcal{C}$ is mainly involved in attribute identification which follows the same rule as used when evaluating projection models’ performance in Sec. 4.1. The complete inference process is also demonstrated in Fig. 8 in Appendix.
@@
5 Ablation Study
We discover that using a pretrained concept space with learned abstract knowledge helps modality-specific projection models converge faster compared to the ones without the access. Specifically, we cut our framework’s access to the pretrained concept space $\mathcal{C}$. Instead, the framework is only provided with a freshly initialized concept space $\mathcal{C}^{\prime}$ and the loss function during pretraining of the vision-modality projection model is changed to $\mathcal{L}_{\text{vision}}^{\prime}=\mathcal{L}_{\text{vision}}+\mathcal{L}_{concept}$. Fig. 3 shows that the original framework’s projection models can converge faster than the ablated version. Based on this evidence, we conclude that the abstract knowledge shared by the pretrained concept space streamlines the learning process of modality-specific projection models.
@@ 6 Discussion
Addressing Bias. Hidden bias learned from datasets often hinders the trustworthiness of ML systems. For example, NLP models often tend to associate the word "monarch" more with the word "male" than "female," reflected, for instance, in higher similarity scores for embeddings of "monarch" and "male." Our proposed framework facilitates effective probing into the model’s learned knowledge and offers the capacity to rectify such biases. Further demonstrations of probing into the learned concept space can be found at Appendix A.3. In the same monarch example, as training targets for concept space are simply probability distributions, bias can be easily addressed by ensuring the ground truth concept relations reflect the same entailment probability between the concept pairs of "monarch-male" and "monarch-female," which could be easily achieved from user interference.
Scalability of the Concept Space. In our experiments, the concept space is organized to reflect ground truth entailment probabilities observed in the training sets. We believe that our approach of replicating entailment probabilities from training sets can be extended to datasets with a broader array of concepts. Previous works (Vilnis et al., 2018; Li et al., 2018; Lai & Hockenmaier, 2017) have demonstrated that similar embedding spaces can accurately learn entailment probabilities for concept pairs in WordNet (WordNet, ). Scaling up the number of concepts introduces a challenge in generating the ground truth of entailment probabilities. We think the rich textual data available today offers a viable avenue for extracting concept relations, including entailment relations, as shown in the work by He & Peng (2020). To further verify the scalability of the concept space, we used the proposed method and fiited an concept space to the full WordNet noun entries, contributing to 10765 concepts. Measured by the KL divergence metric, the WordNet’s concept space achieves a $D_{KL}$ of $0.1308$ against ground truth. For comparison, GQA’s concept space is measured at $0.1172$.
Call for Concept-Focused Datasets. In our development, we discovered a lack of high-quality datasets focused on annotating concepts in real-life images. Even with our preprocessing steps, the attribute/concept annotations in COCO and GQA are significantly noisy, partially reflected in the reduced performance of both our framework and others. We believe that potential datasets with accurate concept annotations could not only benefit the learning of our framework but also aid in the development of more reliable and safer AI systems.
Future Works. Although our results are encouraging, we believe there is room for improvement. The current framework supports a moderate number of concepts defined by entailment relations. We envision future iterations expanding this capability to support more concepts with diverse relations. The results of the Image-Text Matching Task inspire us to explore the potential adaptation of the proposed framework to the Text-to-Image Generation task (Ramesh et al., 2022). The concept space embedded with interpretable knowledge could contribute to achieving a safer and bias-free generative process.
Conclusion. In this work, we introduce a novel multi-modality framework that centers around a concept space embedded with modality-agnostic knowledge. Our experiments show this concept-centric framework demonstrates more efficient learning curves compared to traditional architectures while maintain comparable performances on downstream tasks.
@@ A.2 Concept Space Training Objective
[…]
@@ A.3 Probing into Concept Space
Concept 1 Concept 2 Concept Space Ground Truth
Orange Bus 0.043 0.043
Old Building 0.032 0.048
Smiling Person 0.074 0.073
White Snow 0.910 0.974
Parked Car 0.228 0.244
Cloudy Sky 0.173 0.192
Table 4: Sample Queries of Concepts in GQA
Table 4 shows another example of probing learned knowledge from a concept space fitted to a dataset with a greater array of concepts. Our framework enables easy querying of targeted concept pairs, which would be computationally expensive if not infeasible in traditional latent spaces.
Appendix B Evaluation Datasets and Preprocessing
@@
We base our evaluations on three datasets:
@@
CLEVR dataset comprises synthesized images paired with intricate questions testing a system’s visual reasoning capabilities. We choose CLEVR for evaluation because it provides a highly controlled mini-world, where concepts are easily drawn from visual objects, and relationships between concepts are clearly defined. Each CLEVR image displays a scene with a random number of objects, each described by color, shape, material, and size, which produces 15 unique values, such as blue, cube, forming attribute concepts related to specific objects.
@@
COCO dataset exposes our framework to a knowledge world resembling the real world better than computer-generated images from CLEVR. We use attribute annotations proposed by Patterson & Hays to establish attribute concepts such as soft, cooked, and parked (2016). The original COCO classes are used as category concepts. We focus our evaluation on the top 35 frequent attributes and their associated categories to gain meaningful insights, resulting in 64 concepts.
@@
GQA dataset is similar to COCO, providing a controlled sandbox mimicking the real world. We use the original attribute and category labels in GQA as concepts and filter out rare attributes and classes, resulting in the same amount of concepts as in COCO. Example attribute and category concepts include happy, old, gray, and boy.

### v2

@@ 4.4 Downstream Tasks
4.4.1 Image-Text Matching
Image-text matching is a binary classification task on whether a natural language sentence describes an image. Our framework can naturally adopt a common approach involving creating representations for sentences and images in a shared latent space. In contrast to those works, however, our latent space is a knowledge-embedded concept space that supports efficient probing. Specifically, given an image-text pair $(x_{m}^{\text{vision}},x_{n}^{\text{NL}})$, their representations in the learned concept space $\mathcal{C}$ are generated by $f_{\text{vision}}(x_{m}^{\text{vision}})=\Omega_{m}^{\text{vision}}$ and $f_{\text{NL}}(x_{n}^{\text{NL}})=\Omega_{n}^{\text{NL}}$. The probability that $(x_{m}^{\text{vision}},x_{n}^{\text{NL}})$ is a positive pair can be determined by the cross entailment probability as follows:
$P(\text{matched}\mid(x_{m}^{\text{vision}},x_{n}^{\text{NL}}))=\frac{1}{2}\left[P(\Omega_{m}^{\text{vision}}\mid\Omega_{n}^{\text{NL}})+P(\Omega_{n}^{\text{NL}}\mid\Omega_{m}^{\text{vision}})\right]$
This inference process is demonstrated in Fig. 8 in Appendix.
@@ 4.4.1 Image-Text Matching
To compare our framework’s performance, we implement other benchmark multi-modality models with applications in the Image-Text Matching task. The results are summarized in Table 4. In contrast to those models with traditional black-box architectures, our framework displays a more efficient learning process and adopts a more transparent inference process without sacrificing its performance. Details of this experiment can be found in Appendix E.
@@
4.4.2 Visual Question Answering
@@
Inference Process. A correctly predicted program sequence $\hat{z_{i}}$ starts with a Scene function that returns all objects in an image and ends with a program that outputs the answer $o_{i}$. Intermediate programs take the output from previous programs as inputs, which is a recurring process until the final function. Our concept space $\mathcal{C}$ is mainly involved in attribute identification, following the same procedure used when evaluating projection models in Sec. 4.1. Spatial information for each object is obtained from the bounding-box coordinates predicted by $f_{\text{detection}}$ and is passed to the corresponding symbolic functions that handle spatial reasoning. The complete inference process is also demonstrated in Fig. 9 in Appendix.
@@
Table 5: A comparison between our framework’s performance and state-of-the-art models.
@@ 5 Discussion
A Cognition-Inspired Learning Paradigm. Most current multi-modality learning frameworks, and even the broader landscape of machine learning systems, rely on a learning paradigm that differs substantially from those observed in human cognition. When exposed to new knowledge, we instinctively form a concept and associate it with the external stimuli tied to that information. This newly formed concept is then integrated into our existing body of knowledge and stored as persistent memory in the mind. In contrast, most machine learning frameworks encode knowledge into large sets of model parameters that are difficult to interpret without specific model input. As a result, the activation of learned knowledge in such systems is often transient and dependent on specific inputs. This fundamental difference presents a significant challenge in designing systems that can explicitly form, retain, and reason over interpretable concepts in a manner analogous to human cognition.
The inclusion of a structured concept space and the use of concept-grounded inference may initially appear restrictive, particularly when compared with conventional models that rely on dense, task-specific representations optimized end to end for performance. However, we view this design as a deliberate and principled choice. By introducing a concept space that reflects structured, abstract knowledge similar to how humans form and retain concepts, the framework gains several benefits that are otherwise difficult to achieve. These include more efficient learning, natural generalization across modalities, and interpretability through explicit probing. The concept space acts as an inductive bias consistent with human cognition, enabling machine learning systems to operate in a more principled and cognitively grounded manner. We believe this work highlights a compelling direction for rethinking learning systems to more closely mirror human intelligence.
@@
Addressing Bias. Hidden biases learned from datasets often hinder the trustworthiness of ML systems (Amodei et al., 2016; Lederer, 2023; Kaur et al., 2022; Knott et al., 2023). For example, NLP models often tend to associate the word “monarch” more with the word “male” than “female,” reflected, for instance, in higher similarity scores between embeddings of “monarch” and “male.” Our proposed framework facilitates effective probing into the model’s learned knowledge and offers the capacity to rectify such biases.
Concept 1 Concept 2 Concept Space Ground Truth
Orange Bus 0.043 0.043
Old Building 0.032 0.048
Smiling Person 0.074 0.073
White Snow 0.910 0.974
Parked Car 0.228 0.244
Cloudy Sky 0.173 0.192
Table 6: Sample Entailment Relation Queries of Concepts in Learned GQA Concept Space
Table 6 shows probing of a learned concept space fitted to the GQA dataset in action. Our framework enables easy querying of targeted concept pairs, which would be computationally expensive, if not infeasible, in traditional latent spaces. Further demonstrations of probing into the learned concept space can be found in Appendix A.3.
Revisiting the earlier example of the concept pair “monarch” and gender, the bias can be addressed directly in our framework by adjusting the ground-truth entailment probabilities. Specifically, ensuring equal entailment probabilities between “monarch–male” and “monarch–female” mitigates representational bias, a correction that can be easily applied through user-guided specification.
Scalability of the Concept Space. In our experiments, the concept space is constructed to reflect ground-truth entailment probabilities observed in training data. This approach can scale to larger and more diverse sets of concepts. Prior work (Vilnis et al., 2018; Li et al., 2018; Lai and Hockenmaier, 2017) has shown that similar embedding structures can learn entailment relations for large ontologies such as WordNet (WordNet, ). Although scaling introduces challenges in generating ground-truth probabilities, textual corpora offer a promising resource for extracting such relations, as demonstrated by He and Peng (2020). To assess scalability, we fitted a concept space to the full set of WordNet noun entries, totaling 10,765 concepts. The resulting space achieved a KL divergence of 0.1308 with respect to the ground truth, compared to 0.1172 for the GQA concept space.
Call for Concept-Focused Datasets. A major bottleneck in concept-centric learning is the lack of high-quality datasets with accurate concept annotations. In our experience, even after preprocessing, concept and attribute labels in datasets such as COCO and GQA contain substantial noise. This limits not only the performance of our framework but also that of other systems. Recent works (Pham et al., 2021; Saini et al., 2022; Bravo et al., 2023; Vedantam et al., 2020) have begun addressing this gap by improving the collection and annotation of attribute-focused datasets. We believe that further efforts to build datasets with richer and more reliable concept annotations will greatly support the development of interpretable and trustworthy AI systems.
Integrating Concept Spaces with Vision-Language Models. Compared to conventional vision-language foundation models, our cognitively inspired framework maintains an explicit and interpretable concept space for retaining and reasoning with knowledge, but its reasoning capabilities are currently limited to more controlled and simpler task domains. In contrast, modern VLMs, while operating largely as black boxes, offer strong flexibility and broad generalization across diverse downstream tasks through large-scale pretraining. A promising direction for future work is to explore hybrid architectures that combine these strengths. A concept space, which could take a form similar to that adopted here or extend beyond capturing only entailment relations, could be incorporated into a transformer-based VLM as a set of learned model parameters. Attention mechanisms can then be used to associate model inputs with the knowledge embedded in the concept space. These concept parameters could be kept frozen, allowing the model to use them as a stable semantic memory, or updated during downstream training, enabling the model to refine or even discover new concepts through interaction with task data.
Scope and Limitations. In the current framework, the concept space supports only two types of concepts: attributes and categories. Future work should explore methods to expand this space to a broader range of concepts, such as abstract or relational concepts for more complex reasoning, and action-oriented concepts to accommodate modalities beyond vision and language, such as robot planning.
6 Conclusion
Human cognition exhibits a remarkable ability to form a coherent and structured understanding of the world, and to apply this knowledge efficiently across diverse tasks and modalities. Inspired by this capability, we propose a concept-centric multi-modality learning framework centered around a modality-agnostic concept space that captures universally applicable knowledge.
The primary technical contribution lies in the design of a modular framework that integrates a shared concept space with a flexible set of modality-specific projection models. This design enables knowledge reuse across modalities and task domains, supporting learning that is interpretable, generalizable, and modular. Unlike traditional end-to-end learning systems that encode knowledge implicitly within dense parameter spaces, the proposed framework embeds knowledge explicitly into a structured concept embedding space, enabling interpretability through efficient probing of concept entailment probabilities.
Empirically, the experiments demonstrate that the proposed framework supports more efficient learning. In the vision modality, the projection model converges significantly faster than a baseline model built on a traditional architecture. This gain in efficiency is attributed to the fact that the concept space already encodes structured, abstract knowledge that the projection model can adapt to. Additionally, we evaluated the framework on two downstream tasks, Image-Text Matching and Visual Question Answering, and demonstrated that our method achieves performance comparable to state-of-the-art methods even without task-specific fine-tuning. While our goal is not to surpass existing benchmarks in raw performance, these results support the viability of a cognitively inspired learning paradigm. Rather than optimizing solely for accuracy, our framework emphasizes learning efficiency, interpretability, and structural alignment with human cognition. These qualities are increasingly important as machine learning systems are deployed in more complex and dynamic environments.
More broadly, the work motivates a rethinking of how machine learning systems acquire and represent knowledge. By introducing a concept space as an inductive bias, this framework opens a promising research direction toward building systems that more naturally align with human reasoning. Such systems may offer greater transparency, flexibility, and the ability to generalize knowledge across tasks and modalities.
Looking forward, several directions can further extend the proposed framework. First, scaling the concept space to support larger vocabularies and richer relational structures beyond entailment, including compositional and causal relations, would expand its expressive power. Second, applying the framework to new task domains such as concept-grounded Text-to-Image generation presents a natural extension. Beyond these extensions, the results point toward meta-learning approaches that enable learning systems to discover, update, and organize concepts directly from multimodal inputs, including vision and language. Rather than treating concept vocabularies as fixed and externally provided, future systems should be able to infer and refine conceptual structures from raw data as models interact with new modalities and tasks. Advancing adaptive concept learning mechanisms for concept discovery, concept organization, and concept application is an important step toward more autonomous, scalable, and cognitively grounded machine learning systems. Finally, explicit concept-level representations may serve as a semantic interface for semantic communication, in which agents exchange meaning at the level of concepts and relations rather than raw signals or low-level features.
Acknowledgements
This work was supported in part by the National Science Foundation under Grant No. 2133403, and by computing credits provided by Amazon Web Services and Microsoft Azure. We thank the reviewers and the action editor for their thoughtful feedback and constructive suggestions.
@@ A.2 Concept Space Training Objective
$\begin{split}\mathcal{L}_{\text{concept}}(\mathcal{C};&\mathcal{D}_{*})=\frac{1}{|\mathcal{D_{*}}|}\sum_{(x,\bm{y})\in\mathcal{D}_{*}}\\ &\frac{1}{2\cdot\mathinner{\!\left\lvert\binom{\bm{y}}{2}\right\rvert}}\sum_{(y_{1},y_{2})\in\binom{\bm{y}}{2}}D_{\mathbf{KL}}(P(y_{1}|y_{2})||Q(y_{1}|y_{2}))+D_{\mathbf{KL}}(1-P(y_{1}|y_{2})||1-Q(y_{1}|y_{2}))\end{split}$ (6)
@@ A.3 Probing into Concept Space
A.4 Ablation on Concept Space
We discover that using a pretrained concept space with learned abstract knowledge helps modality-specific projection models converge faster compared to the ones without the access. Specifically, we cut our framework’s access to the pretrained concept space $\mathcal{C}$. Instead, the framework is only provided with a freshly initialized concept space $\mathcal{C}^{\prime}$ and the loss function during pretraining of the vision-modality projection model is changed to $\mathcal{L}_{\text{vision}}^{\prime}=\mathcal{L}_{\text{vision}}+\mathcal{L}_{\mathcal{C}}$. Fig. 6 shows that the original framework’s projection models can converge faster than the ablated version. Based on this evidence, we conclude that the abstract knowledge shared by the pretrained concept space streamlines the learning process of modality-specific projection models.
A.5 Ablation on Concept Space Dimensionality
Dataset Dim ($K$) Accuracy mAP
CLEVR 24 0.99898
50 0.99900
96 0.92914
COCO 24 0.95899 0.57212
50 0.95653 0.58961
96 0.56060 0.56024
GQA 24 0.84752 0.21271
50 0.84419 0.35202
96 0.61458 0.32331
Table 7: Ablation on concept space dimensionality ($K$). CLEVR reports accuracy only; COCO and GQA report accuracy and mAP. Bolded values indicate the selected dimensionality ($K=50$).
To assess how the dimensionality of the concept space affects downstream performance, we conduct an ablation study using $K\in\{24,50,96\}$. For each value of $K$, we follow the same pretraining protocol: we first train the concept space to convergence and then train the ViT-based projection model to adapt to that space. All runs use identical hyperparameters, optimization settings, and training steps to ensure a fair comparison. Results are reported in Table 7.
Empirically, $K=50$ achieves the strongest and most consistent performance across datasets, showing the highest mAP on both COCO and GQA while matching the near-saturated accuracy on CLEVR. These observations provide practical support for choosing $K=50$ as the concept-space dimensionality used in all main experiments.
We also observe that increasing the dimensionality to $K=96$ leads to noticeable performance degradation, particularly on COCO and GQA. A plausible explanation is that larger box dimensions introduce a more complex geometric structure, which makes the concept space harder to learn effectively and can lead to diminishing returns or overfitting. Conversely, very small dimensions, such as $K=24$, limit expressivity and may restrict performance on semantically richer datasets. The intermediate dimensionality of $K=50$ therefore offers a favorable balance between expressiveness and learnability.
Appendix B Cross Modality Joint Training
To allow probabilistic analysis for cross-modality tasks, an optional joint training stage can be used to encourage different projection models to produce projections that overlap with each other’s for the same object. This joint training stage is optional because each individual projection model is already adapted to the shared knowledge space, and it is computationally lightweight, as the modality-specific projection models have already been trained and aligned with the unified concept space. It requires very modest resources, with convergence occurring within a few hundred training steps, as demonstrated in Sec. 4.2. Subsequently, this design with demonstrated efficiency allows the effortless incorporation of new projection models into our proposed framework, mirroring humans’ ability to learn and link knowledge across modalities in a fast and efficient manner. Specifically, consider a system with two modalities, A and B, as an example. The training dataset would be denoted as $\mathcal{D}_{A\cup B}=\{(x^{A}_{i},x^{B}_{i},\bm{y}_{i})\}_{i=1}^{N}$, and the training objective for this joint training stage is defined as:
$\begin{split}\mathcal{L}&{}_{\text{joint}}(\theta_{A},\theta_{B};\mathcal{D}_{A\cup B})=\frac{1}{2|\mathcal{D}_{A\cup B}|}\sum_{(x_{A},x_{B},\bm{y})\in\mathcal{D}_{A\cup B}}\\ &P(f_{A}(x_{A};\theta_{A})\mid f_{B}(x_{B};\theta_{B}))+P(f_{B}(x_{B};\theta_{B})\mid f_{A}(x_{A};\theta_{A}))\end{split}$ (7)
The overall training objective becomes a combination of modality-specific projection losses and this joint training loss.
Appendix C Evaluation Datasets and Preprocessing
C.1 Dataset Details
Dataset # Attributes # Categories
GQA 35 33
CLEVR 12 3
COCO 35 29
Unified (All) 70 60
Table 8: Summary of concept subsets used in each dataset.
Type CLEVR COCO GQA
Attributes blue, brown, cyan, gray, green, large, metal, purple, red, rubber, small, yellow adult, appetizing, athletic, busy, casual, cloth, cooked, enjoying, family-friendly, female, fluffy, fresh, functional, furry, hairy, healthy, holding, horizontal, laying, male, metal/metallic, moving, parked, participating, public, sitting, socializing, soft, sporty, standing, tame, tasty/delicious, useful, vertical, watching/looking black, blue, brick, brown, clear, cloudy, concrete, dark, glass, gray, green, happy, large, long, metal, old, open, orange, parked, pink, red, round, short, silver, sitting, small, smiling, standing, striped, tall, tan, white, wood, yellow, young
Categories cube, cylinder, sphere airplane, apple, banana, bear, bicycle, bird, boat, broccoli, bus, cake, car, carrot, cat, cow, dog, donut, elephant, giraffe, horse, hot dog, motorcycle, orange, person, pizza, sandwich, sheep, train, truck, zebra bed, boy, building, bus, car, chair, fence, field, floor, giraffe, girl, grass, ground, hair, head, jacket, man, person, plate, road, shirt, sidewalk, sky, snow, street, table, train, tree, trees, wall, water, window, woman
Table 9: Exact attribute and category concepts present in each dataset.
Concept Type Datasets
blue attribute CLEVR, GQA
brown attribute CLEVR, GQA
gray attribute CLEVR, GQA
green attribute CLEVR, GQA
large attribute CLEVR, GQA
metal attribute CLEVR, GQA
parked attribute COCO, GQA
red attribute CLEVR, GQA
sitting attribute COCO, GQA
small attribute CLEVR, GQA
standing attribute COCO, GQA
yellow attribute CLEVR, GQA
bus category COCO, GQA
car category COCO, GQA
giraffe category COCO, GQA
person category COCO, GQA
train category COCO, GQA
Table 10: Overlapping concepts appearing in multiple datasets. These concepts are merged in the Unified dataset.
@@ C.1 Dataset Details
We base our evaluations on three datasets and a world dataset aggregating all concepts and representations together:
@@
CLEVR dataset comprises synthesized images paired with intricate questions testing a system’s visual reasoning capabilities. We choose CLEVR for evaluation because it provides a highly controlled mini-world, where concepts are easily drawn from visual objects, and relationships between concepts are clearly defined. Each CLEVR image displays a scene with a random number of objects, each described by color, shape, material, and size, which produces 15 unique values such as blue and cube, forming attribute concepts related to specific objects.
@@
COCO dataset exposes our framework to a knowledge world resembling the real world better than computer-generated images from CLEVR. We use attribute annotations proposed by Patterson & Hays to establish attribute concepts such as soft, cooked, and parked (see Fig. 1 in (2016)). The original COCO object categories are used as category concepts. We focus our evaluation on the top 35 frequent attributes and their associated categories to gain meaningful insights, resulting in 64 concepts (see Table 9).
@@
GQA dataset is similar to COCO, providing a controlled sandbox mimicking real-world settings. We use the original attribute and category labels in GQA as concepts and filter out rare attributes and classes, resulting in the same total number of concepts as COCO. Example attribute and category concepts include happy, old, gray, and boy. Exact concepts can be found in Table 9.
Unified dataset aggregates all attribute and category concepts appearing in CLEVR, COCO, and GQA into a single world dataset, resulting in 130 unique concepts (70 attributes and 60 categories; see Table 8). Overlapping concepts across datasets (listed in Table 10) are merged into unified concepts. This dataset allows us to evaluate the scalability of the proposed framework and its ability to generalize abstract concepts across different visual and linguistic domains. To avoid the Unified dataset being dominated by CLEVR, which contains significantly more samples than COCO and GQA, we balance the dataset sizes during construction. In the raw training splits, CLEVR contains 455,632 samples, compared to 78,898 in GQA and 91,667 in COCO. We therefore cap the CLEVR training split to match the size of the largest non-CLEVR dataset (COCO), resulting in 91,667 CLEVR samples after filtering. This ensures that the Unified dataset remains balanced and diverse across data sources.
C.2 Dataset Preprocessing